# Hardware-Assisted Virtualization of Neural Processing Units for Cloud Platforms

### v1

@@ IV Discussion
Support for multi-chip inferences. Currently, TCloud supports multi-chip inference with data parallelism by using multiple vNPU chips. As the first step of NPU virtualization, we focus on enabling fine-grained resource sharing on individual NPU chips. In future work, we will extend TCloud by investigating how to virtualize inter-chip interconnects to support more complicated scenarios (e.g., model parallelism).
Engineering efforts in developing TCloud. While TCloud is a full-stack NPU virtualization design, each component is developed and tested in a modular way, and we minimize changes to the existing system at each level to reduce the debugging and verification efforts. The compiler and hardware changes are minimized as NeuISA reuses the VLIW instruction format in each $\mu$TOp. The guest vNPU driver greatly resembles a native NPU driver thanks to PCIe pass-through, and the major change is the new hypercalls for vNPU management. The KVM hypervisor already provides extensibility for new PCIe devices, and we leverage this feature to integrate the vNPU manager.
Inter-generational compatibility with NeuISA. NeuISA enables a DNN program to run on different numbers of MEs/VEs without recompilation. This greatly eases the effort to provide compatibility across generations of NPU hardware. NeuISA could ease the future development efforts of ML frameworks to support new NPU hardware and enable more flexible and transparent ways to manage NPU resources. TCloud provides a general vNPU abstraction, which allows a vNPU to be mapped to different generations of NPU hardware.
@@ V Evaluation
Our evaluation shows that: (1) TCloud provides performance isolation with up to 4.6×4.6$\times$4.6 × reduction in tail latency, while improving the ML service throughput by 1.4$\times$ over state-of-the-art NPU sharing approaches (§$\S$§V-B); (2) It improves the NPU utilization by 1.2$\times$ (§$\S$§V-C); (3) It scales as we change the number of MEs/VEs (§$\S$§V-E); (4) It benefits multi-tenant ML services with various HBM bandwidths (§$\S$§V-F).
@@ V-A Experimental Setup
We evaluate DNN workloads (see Table I) from MLPerf v2.1 [48] and the official TPU reference models [23]. To test TCloud under different workload combinations, we select workload pairs with low ME/VE contention (DLRM+SMask, DLRM+RtNt, NCF+RsNt), medium contention (ENet+SMask, BERT+ENet, ENet+MRCN), and high contention (ENet+TFMR, MNIST+RtNt, RNRS+RtNt). The batch size is 32 except for MRCN and SMask (batch size is 8 for them). Each workload runs on a vNPU with 2 MEs and 2 VEs. We map two vNPUs to a physical NPU core with 4 MEs and 4 VEs as listed in Table II. The SRAM and HBM capacity is evenly partitioned between the vNPUs. To obtain steady-state performance, we run inference requests continuously for each workload until all collocated workloads have completed a certain number of requests. We compare the following designs:
@@
  • TCloud-NoHarvest (TCloud-NH): spatial-isolated vNPUs with dedicated MEs/VEs without dynamic scheduling. This resembles existing static partitioning techniques such as NVIDIA Multi-instance GPU (MIG) [43].
  • TCloud: spatial-isolated vNPUs with dynamic resource scheduling and harvesting enabled by NeuISA.
Figure 19: 95% Percentile latency of TCloud (normalized to PMT).
Figure 20: Average request latency of TCloud (normalized to PMT).
Figure 21: Throughput of TCloud (normalized to PMT).
@@
V-B Performance of TCloud
Tail Latency. Figure 19 shows that TCloud improves the 95% tail latency over V10 by up to 4.6×4.6$\times$4.6 × (1.56×1.56$\times$1.56 × on average). V10 primarily focuses on maximizing the utilization. Thus, even with an operator preemption mechanism, it still fails to enforce performance isolation between vNPUs, due to complex inter-operator dependencies and imbalanced operator lengths.
In contrast, TCloud ensures performance isolation between vNPUs while opportunistically improving their performance by harvesting. As TCloud only harvests the underutilized compute units, the performance interference between vNPUs is minimized. Hence, a harvested workload in TCloud experiences negligible tail latency overhead compared to that in TCloud-NH. In a few cases (e.g., ENet+MRCN and RNRS+RtNt), harvesting increases the burden on memory bandwidth, which may slightly impact the tail latency. However, TCloud still achieves much better tail latency than PMT and V10.
Average Latency. TCloud improves the average latency of inference requests by 1.33$\times$ over PMT and 1.12$\times$ over V10 on average (Figure 20). While both V10 and TCloud perform dynamic scheduling to utilize the NPU hardware, TCloud greatly reduces ME contentions with $\mu$TOp-level scheduling. V10 treats all MEs on a physical NPU core as a whole unit, due to the VLIW ISA limitation. This causes false contentions on the MEs when an operator cannot fully exploit the MEs but still fully occupies them. In contrast, TCloud eliminates such contention by assigning MEs to different operators.
Figure 23: Benefit breakdown of ME/VE harvesting. Y-axis is the speedup of TCloud over TCloud-NH. X-axis is the percentage of time when executing on TCloud-NH. The curves below $Y=1$ indicate the slowdown of operators due to interference. The curves above $Y=1$ indicate the speedup due to harvesting.
@@ V-B Performance of TCloud
Throughput. Figure 21 shows the throughput of the collocated workloads. When the ME/VE contention is low, both V10 and TCloud improve the throughput significantly over PMT (by 1.58×1.58$\times$1.58 × and 1.62×1.62$\times$1.62 × on average), as the major benefit comes from overlapping the execution of ME-intensive operators and VE-intensive operators. When the ME/VE contention is high, TCloud improves the throughput of DNN workloads over V10 by up to 1.41$\times$, since TCloud offers more flexibility for dynamic ME/VE scheduling with $\mu$TOps, as discussed above.
@@ V-C Resource Utilization Improvement
We show the utilization of the MEs and VEs on the NPU core in Figure 22. With dynamic operator scheduling, TCloud improves the ME and VE utilization by 1.26$\times$ and 1.2$\times$ over PMT on average. For some workload pairs, TCloud achieves slightly better utilization than V10, since TCloud has less preemption overhead with $\mu$TOp scheduling. Specifically, V10 needs to preempt the entire operator from all MEs, while TCloud only preempts the $\mu$TOps on the harvested MEs, such that the remaining $\mu$TOps can continue execution.
@@ V-D Benefit Breakdown of ME/VE Harvesting
To better understand the benefit and overhead of harvesting, we trace the speedup of each operator in TCloud over TCloud-NH, and show the impact in Figure 23.
@@
To visualize the behavior of TCloud’s dynamic ME/VE scheduling, we trace the number of MEs and VEs assigned to each collocated workload at runtime in Figure 24. As the ME/VE demands of the workloads vary across time, the ME-intensive workload (e.g., RetinaNet and ShapeMask) attempts to harvest the unused MEs from the collocated workload. The VEs are harvested similarly.
@@ V-E Impact of Varying MEs and VEs
Figure 25: Throughput improvement of TCloud with varying numbers of MEs and VEs over V10 with 2 MEs and 2 VEs.
To show TCloud’s benefits on different hardware configurations, we vary the numbers of MEs and VEs on the physical NPU core and evenly partition the core between the two collocated vNPUs. We compare TCloud with V10, as V10 has fine-grained preemption, which serves as the most competitive baseline. We show the throughput in Figure 25. With more MEs/VEs, TCloud brings more benefits, since there is more flexibility for dynamic ME/VE scheduling. With more MEs/VEs, it is more likely that an operator cannot fully exploit all ME/VEs. Therefore, the benefit of $\mu$TOp-level scheduling and harvesting becomes more obvious.
@@ V-F Impact of Varying Memory Bandwidth
Figure 26: Throughput improvement of TCloud with varying HBM bandwidth (normalized to V10).
We show TCloud’s performance under different HBM bandwidth configurations in Figure 26. In most cases, TCloud achieves similar throughput benefits. This is because many ML inference workloads suffer from the ME/VE contention rather than HBM bandwidth contention (see Figure 7). To understand the impact of memory bandwidth contention, we collocate two memory-intensive workloads (i.e., DLRM+NCF and NCF+TFMR). Even with low available memory bandwidth (e.g., 900 GB/s), TCloud still outperforms the time sharing-based scheme V10. With more available bandwidth, TCloud brings more benefits for memory-intensive workloads, since higher bandwidth helps alleviate memory contention.
For memory-intensive workloads, TCloud enables them to be collocated with compute-intensive workloads following existing workload collocation approaches [57, 12, 39], which helps cloud platforms better utilize both compute and memory resources. As a case study, we collocate a memory bandwidth-intensive LLM inference workload, LLaMA2-13B [52] (LLaMA), with compute-intensive workloads (i.e., BERT, RsNt, and RtNt). As shown in Figure 27, with V10, when LLaMA temporarily occupies all MEs/VEs, it underutilizes the MEs/VEs since the execution is bounded by memory bandwidth. However, the underutilized MEs cannot be harvested by the collocated workload due to the temporal sharing mechanism. TCloud enables the spatial sharing of the MEs/VEs, so the collocated workload can harvest the spare MEs/VEs for throughput improvement (by up to 1.6$\times$). Meanwhile, LLaMA suffers from negligible overhead while using fewer MEs/VEs. As hardware vendors continue to scale the HBM bandwidth, TCloud will bring more benefits because of the alleviation of memory contention. Note that the owners of LLM inference service have the choice to follow the pay-as-you-go model to allocate multiple vNPUs with large memory, the vNPU abstraction of TCloud offers the flexibility for resource allocation while enabling dynamic scheduling.
Figure 27: Performance of collocating LLM (LLaMA-13B with batch size 8 and input sequence length 512) and other models in TCloud.
@@ VI Related Work
System virtualization for accelerators. As we employ hardware accelerators for ML services, cloud platforms prefer to virtualize them for improved resource utilization [6, 56, 27, 61, 35, 62, 34]. Prior studies have investigated virtualization techniques for GPUs [44, 43, 51, 37] and FPGAs [34, 40, 61, 35, 62]. Unfortunately, these techniques cannot be directly applied to NPUs, as they target different architectures. AvA [60] investigates hypervisor interposition techniques for virtualizing accelerators. However, they do not focus on improving the resource utilization. To the best of our knowledge, TCloud is the first to investigate the system and architectural techniques for NPU virtualization.
While different tensor processors have been developed recently [36, 31, 8, 49], most of them have specialized matrix engines and generic vector engines, given the continuing trend that DNN computations are dominated by these operations. As the imbalanced ME/VE demands are intrinsic to DNN workloads (§$\S$§II-B), these processors also suffer from resource underutilization. The design of TCloud can be adapted to virtualize these accelerators for utilization improvement.
Accelerator resource sharing and scheduling. There have been various techniques [27, 42, 13, 12, 34, 40, 61, 35, 62] for supporting multi-tenant workloads on accelerators. PREMA [16] proposed a preemptive scheduling mechanism, but it causes high context-switch overhead. Prior studies [9, 45, 57] investigated the imbalance of compute units and memory. Planaria [20] studied the spatial underutilization of systolic arrays. V10 [57] enabled fine-grained preemption. There are also software techniques to fuse DNN workloads at graph level [59, 58]. However, they force two DNN inference tasks to launch together, which cannot work for the unpredictable incoming requests in the cloud. None of them systematically enables NPU virtualization. TCloud addressed the systems and architectural challenges of NPU virtualization.
As the resource demand of a DNN workload changes drastically over time (see §II-B), a static resource allocation is insufficient. TCloud proposes NeuISA and enables dynamic scheduling to mitigate the underutilization, it is orthogonal to the higher-level workload collocation techniques.
@@ VII Conclusion
We identify the key challenges of virtualizing NPUs for cloud platforms, including the need for fine-grained system abstraction and resource scheduling and the necessity of architectural support. We present a holistic solution TCloud for enabling NPU virtualization. It improves both NPU utilization and performance isolation for multi-tenant ML services.

### v2

@@ IV Discussion
Support for multi-chip inferences. Currently, Neu10 supports multi-chip inference with data parallelism by using multiple vNPU chips. As the first step of NPU virtualization, we focus on enabling fine-grained resource sharing on individual NPU chips. In future work, we will extend Neu10 by investigating how to virtualize inter-chip interconnects to support more complicated scenarios (e.g., model parallelism).
Engineering efforts in developing Neu10. While Neu10 is a full-stack NPU virtualization design, each component is developed and tested in a modular way, and we minimize changes to the existing system at each level to reduce the debugging and verification efforts. The compiler and hardware changes are minimized as NeuISA reuses the VLIW instruction format in each $\mu$TOp. The guest vNPU driver greatly resembles a native NPU driver thanks to PCIe pass-through, and the major change is the new hypercalls for vNPU management. The KVM hypervisor already provides extensibility for new PCIe devices, and we leverage this feature to integrate the vNPU manager.
Inter-generational compatibility with NeuISA. NeuISA enables a DNN program to run on different numbers of MEs/VEs without recompilation. This greatly eases the effort to provide compatibility across generations of NPU hardware. NeuISA could ease the future development efforts of ML frameworks to support new NPU hardware and enable more flexible and transparent ways to manage NPU resources. Neu10 provides a general vNPU abstraction, which allows a vNPU to be mapped to different generations of NPU hardware.
@@ V Evaluation
Our evaluation shows that: (1) Neu10 provides performance isolation with up to 4.6×4.6$\times$4.6 × reduction in tail latency, while improving the ML service throughput by 1.4$\times$ over state-of-the-art NPU sharing approaches (§$\S$§V-B); (2) It improves the NPU utilization by 1.2$\times$ (§$\S$§V-C); (3) It scales as we change the number of MEs/VEs (§$\S$§V-E); (4) It benefits multi-tenant ML services with various HBM bandwidths (§$\S$§V-F).
@@ V-A Experimental Setup
We evaluate DNN workloads (see Table I) from MLPerf v2.1 [48] and the official TPU reference models [23]. To test Neu10 under different workload combinations, we select workload pairs with low ME/VE contention (DLRM+SMask, DLRM+RtNt, NCF+RsNt), medium contention (ENet+SMask, BERT+ENet, ENet+MRCN), and high contention (ENet+TFMR, MNIST+RtNt, RNRS+RtNt). The batch size is 32 except for MRCN and SMask (batch size is 8 for them). Each workload runs on a vNPU with 2 MEs and 2 VEs. We map two vNPUs to a physical NPU core with 4 MEs and 4 VEs as listed in Table II. The SRAM and HBM capacity is evenly partitioned between the vNPUs. To obtain steady-state performance, we run inference requests continuously for each workload until all collocated workloads have completed a certain number of requests. We compare the following designs:
@@
  • Neu10-NoHarvest (Neu10-NH): spatial-isolated vNPUs with dedicated MEs/VEs without dynamic scheduling. This resembles existing static partitioning techniques such as NVIDIA Multi-instance GPU (MIG) [43].
  • Neu10: spatial-isolated vNPUs with dynamic resource scheduling and harvesting enabled by NeuISA.
Figure 19: 95% Percentile latency of Neu10 (normalized to PMT).
Figure 20: Average request latency of Neu10 (normalized to PMT).
Figure 21: Throughput of Neu10 (normalized to PMT).
@@
V-B Performance of Neu10
Tail Latency. Figure 19 shows that Neu10 improves the 95% tail latency over V10 by up to 4.6×4.6$\times$4.6 × (1.56×1.56$\times$1.56 × on average). V10 primarily focuses on maximizing the utilization. Thus, even with an operator preemption mechanism, it still fails to enforce performance isolation between vNPUs, due to complex inter-operator dependencies and imbalanced operator lengths.
In contrast, Neu10 ensures performance isolation between vNPUs while opportunistically improving their performance by harvesting. As Neu10 only harvests the underutilized compute units, the performance interference between vNPUs is minimized. Hence, a harvested workload in Neu10 experiences negligible tail latency overhead compared to that in Neu10-NH. In a few cases (e.g., ENet+MRCN and RNRS+RtNt), harvesting increases the burden on memory bandwidth, which may slightly impact the tail latency. However, Neu10 still achieves much better tail latency than PMT and V10.
Average Latency. Neu10 improves the average latency of inference requests by 1.33$\times$ over PMT and 1.12$\times$ over V10 on average (Figure 20). While both V10 and Neu10 perform dynamic scheduling to utilize the NPU hardware, Neu10 greatly reduces ME contentions with $\mu$TOp-level scheduling. V10 treats all MEs on a physical NPU core as a whole unit, due to the VLIW ISA limitation. This causes false contentions on the MEs when an operator cannot fully exploit the MEs but still fully occupies them. In contrast, Neu10 eliminates such contention by assigning MEs to different operators.
Figure 23: Benefit breakdown of ME/VE harvesting. Y-axis is the speedup of Neu10 over Neu10-NH. X-axis is the percentage of time when executing on Neu10-NH. The curves below $Y=1$ indicate the slowdown of operators due to interference. The curves above $Y=1$ indicate the speedup due to harvesting.
@@ V-B Performance of Neu10
Throughput. Figure 21 shows the throughput of the collocated workloads. When the ME/VE contention is low, both V10 and Neu10 improve the throughput significantly over PMT (by 1.58×1.58$\times$1.58 × and 1.62×1.62$\times$1.62 × on average), as the major benefit comes from overlapping the execution of ME-intensive operators and VE-intensive operators. When the ME/VE contention is high, Neu10 improves the throughput of DNN workloads over V10 by up to 1.41$\times$, since Neu10 offers more flexibility for dynamic ME/VE scheduling with $\mu$TOps, as discussed above.
@@ V-C Resource Utilization Improvement
We show the utilization of the MEs and VEs on the NPU core in Figure 22. With dynamic operator scheduling, Neu10 improves the ME and VE utilization by 1.26$\times$ and 1.2$\times$ over PMT on average. For some workload pairs, Neu10 achieves slightly better utilization than V10, since Neu10 has less preemption overhead with $\mu$TOp scheduling. Specifically, V10 needs to preempt the entire operator from all MEs, while Neu10 only preempts the $\mu$TOps on the harvested MEs, such that the remaining $\mu$TOps can continue execution.
@@ V-D Benefit Breakdown of ME/VE Harvesting
To better understand the benefit and overhead of harvesting, we trace the speedup of each operator in Neu10 over Neu10-NH, and show the impact in Figure 23.
@@
To visualize the behavior of Neu10’s dynamic ME/VE scheduling, we trace the number of MEs and VEs assigned to each collocated workload at runtime in Figure 24. As the ME/VE demands of the workloads vary across time, the ME-intensive workload (e.g., RetinaNet and ShapeMask) attempts to harvest the unused MEs from the collocated workload. The VEs are harvested similarly.
@@ V-E Impact of Varying MEs and VEs
Figure 25: Throughput improvement of Neu10 with varying numbers of MEs and VEs over V10 with 2 MEs and 2 VEs.
To show Neu10’s benefits on different hardware configurations, we vary the numbers of MEs and VEs on the physical NPU core and evenly partition the core between the two collocated vNPUs. We compare Neu10 with V10, as V10 has fine-grained preemption, which serves as the most competitive baseline. We show the throughput in Figure 25. With more MEs/VEs, Neu10 brings more benefits, since there is more flexibility for dynamic ME/VE scheduling. With more MEs/VEs, it is more likely that an operator cannot fully exploit all ME/VEs. Therefore, the benefit of $\mu$TOp-level scheduling and harvesting becomes more obvious.
@@ V-F Impact of Varying Memory Bandwidth
Figure 26: Throughput improvement of Neu10 with varying HBM bandwidth (normalized to V10).
We show Neu10’s performance under different HBM bandwidth configurations in Figure 26. In most cases, Neu10 achieves similar throughput benefits. This is because many ML inference workloads suffer from the ME/VE contention rather than HBM bandwidth contention (see Figure 7). To understand the impact of memory bandwidth contention, we collocate two memory-intensive workloads (i.e., DLRM+NCF and NCF+TFMR). Even with low available memory bandwidth (e.g., 900 GB/s), Neu10 still outperforms the time sharing-based scheme V10. With more available bandwidth, Neu10 brings more benefits for memory-intensive workloads, since higher bandwidth helps alleviate memory contention.
For memory-intensive workloads, Neu10 enables them to be collocated with compute-intensive workloads following existing workload collocation approaches [57, 12, 39], which helps cloud platforms better utilize both compute and memory resources. As a case study, we collocate a memory bandwidth-intensive LLM inference workload, LLaMA2-13B [52] (LLaMA), with compute-intensive workloads (i.e., BERT, RsNt, and RtNt). As shown in Figure 27, with V10, when LLaMA temporarily occupies all MEs/VEs, it underutilizes the MEs/VEs since the execution is bounded by memory bandwidth. However, the underutilized MEs cannot be harvested by the collocated workload due to the temporal sharing mechanism. Neu10 enables the spatial sharing of the MEs/VEs, so the collocated workload can harvest the spare MEs/VEs for throughput improvement (by up to 1.6$\times$). Meanwhile, LLaMA suffers from negligible overhead while using fewer MEs/VEs. As hardware vendors continue to scale the HBM bandwidth, Neu10 will bring more benefits because of the alleviation of memory contention. Note that the owners of LLM inference service have the choice to follow the pay-as-you-go model to allocate multiple vNPUs with large memory, the vNPU abstraction of Neu10 offers the flexibility for resource allocation while enabling dynamic scheduling.
Figure 27: Performance of collocating LLM (LLaMA-13B with batch size 8 and input sequence length 512) and other models in Neu10.
@@ VI Related Work
System virtualization for accelerators. As we employ hardware accelerators for ML services, cloud platforms prefer to virtualize them for improved resource utilization [6, 56, 27, 61, 35, 62, 34]. Prior studies have investigated virtualization techniques for GPUs [44, 43, 51, 37] and FPGAs [34, 40, 61, 35, 62]. Unfortunately, these techniques cannot be directly applied to NPUs, as they target different architectures. AvA [60] investigates hypervisor interposition techniques for virtualizing accelerators. However, they do not focus on improving the resource utilization. To the best of our knowledge, Neu10 is the first to investigate the system and architectural techniques for NPU virtualization.
While different tensor processors have been developed recently [36, 31, 8, 49], most of them have specialized matrix engines and generic vector engines, given the continuing trend that DNN computations are dominated by these operations. As the imbalanced ME/VE demands are intrinsic to DNN workloads (§$\S$§II-B), these processors also suffer from resource underutilization. The design of Neu10 can be adapted to virtualize these accelerators for utilization improvement.
Accelerator resource sharing and scheduling. There have been various techniques [27, 42, 13, 12, 34, 40, 61, 35, 62] for supporting multi-tenant workloads on accelerators. PREMA [16] proposed a preemptive scheduling mechanism, but it causes high context-switch overhead. Prior studies [9, 45, 57] investigated the imbalance of compute units and memory. Planaria [20] studied the spatial underutilization of systolic arrays. V10 [57] enabled fine-grained preemption. There are also software techniques to fuse DNN workloads at graph level [59, 58]. However, they force two DNN inference tasks to launch together, which cannot work for the unpredictable incoming requests in the cloud. None of them systematically enables NPU virtualization. Neu10 addressed the systems and architectural challenges of NPU virtualization.
As the resource demand of a DNN workload changes drastically over time (see §II-B), a static resource allocation is insufficient. Neu10 proposes NeuISA and enables dynamic scheduling to mitigate the underutilization, it is orthogonal to the higher-level workload collocation techniques.
@@ VII Conclusion
We identify the key challenges of virtualizing NPUs for cloud platforms, including the need for fine-grained system abstraction and resource scheduling and the necessity of architectural support. We present a holistic solution Neu10 for enabling NPU virtualization. It improves both NPU utilization and performance isolation for multi-tenant ML services.